OpenStudy (anonymous):

Which of the following is NOT equivalent to the rational expression? (Expression Down Below)

OpenStudy (anonymous):

\[\frac{ 9 }{ x-2 }\]

OpenStudy (anonymous):

what are the answer choices?

OpenStudy (anonymous):

\[A: \frac{ -9 }{ 2-x }\] \[B: - \frac{ 9 }{ 2-x }\] \[C: - \frac{ 9 }{ x-2 }\]

OpenStudy (anonymous):

Well, you can set them equal to one another to find out: if the results are equal, the expression is equal and if they aren't, hats you answer.

OpenStudy (anonymous):

??